Supports at A and B are pins. Find a) AD, b) CD, c) DE and d) BC bars. 1 kip = 1000 lb. P kips 12 ft 8 ft 1910 10 ft Sft -
Answer #1

tano = 4 0 = 38-6598 tang=y Ax 12f4 L Vo p. 14ft O ㅏ d u $= 21.80140 Bx = Reaction force at B Axes horizontal Reachon - forceConsider joint F, FOF loos Equating horzontal forces FDF Coso + F-EF COOO=0 FDE = -FEF Equaling vertical forces FDE BIN O - PEquating vertical forces FAD sin 0 = FDF SMO + Fed Bin 0 + FDE FDE = FAD sind - For smo = 0,67314581121.8014 – 0,80039081n38-